A trader gives a discount of 12% on the market price of an article. If a customer pays $1760 for the article. what is the market price of the article
Answer by rothauserc(4718)

let x be the market price
:
1 - 0.12 = 0.88
:
Note the customer paid 88% of the market price
:
0.88 * x = 1760
:
x = 1760/0.88 = 2000
:
************************************
market price of the article is $2000
************************************
: